OpenStudy (anonymous):

solve by completing the sq 3x^2-12x+11=0

OpenStudy (anonymous):

do u know how to solve this problem...... u need to know the procedure

OpenStudy (anonymous):

divide by 3 and take the constant to the other side, then add (b/2)^2 to both sides THat is the procedure

OpenStudy (anonymous):

yes i need step by step please

OpenStudy (anonymous):

\[3x^2-12x=-11\] \[x^2-4x=-\frac{11}{4}\] \[(x-2)^2=-\frac{11}{2}+4=-\frac{3}{2}\] is it ok if the solutions are complex?

OpenStudy (anonymous):

that is wrong satellite

OpenStudy (anonymous):

step one subtract 11 from both sides step 2, divide by 3 yes is certainly made a mistake there!

OpenStudy (anonymous):

no trouble there he is bang on!

OpenStudy (anonymous):

should be \[x^2-4x=-\frac{11}{3}\]

OpenStudy (anonymous):

in second step he is adding 4 on both sides

OpenStudy (anonymous):

then take half the coefficient of the x term (here -4) square it and add to the right to give \[(x-2)^2=-\frac{11}{3}+4=\frac{1}{3}\]

OpenStudy (anonymous):

then take the square root of both sides. you will get two solutions \[x-2=\sqrt{\frac{1}{3}}=\frac{1}{\sqrt{3}}=\frac{\sqrt{3}}{3}\] and \[x-2=-\frac{\sqrt{3}}{3}\]

OpenStudy (anonymous):

finally add 2 to both sides to get your answers. they are \[x=2-\frac{\sqrt{3}}{3}\] and \[x=2+\frac{\sqrt{3}}{3}\]

OpenStudy (anonymous):

thank you you are very helpful:)